What is the steps to Factor this polynomial completely?
2x²-12x-80

Answer by Edwin McCravy(20054)

2x²-12x-80

We factor out common factor 2

2(x²-6x-40) 

We want to factor x²-6x-40 so we
write this:

2(x  )(x  ) 

First we look at the 40

(We work right to left on the x²-6x-40)

Ww write down all the ways two whole numbers
can be multiplied to give 40, they are

40·1
20·2
10·4
 8·5

We look at x²-6x-40 for the sign of the 40. 
It is - so out beside those we SUBTRACT 
those numbers:  [If it were + we would ADD here]

40·1   40-1 = 39
20·2   20-2 = 18 
10·4   10-4 = 6
 8·5    8-5 = 3

Next we look at x²-6x-40 for the 6 and we see that 6 
is listed among those subtractions.  So we use the 
10 and the 4.

We write the 10 in the first parentheses and the 4 in
the second

2(x  10)(x  4)

Next we look at x²-6x-40 for the sign of the 6. It
is -, so we put a - next the larger number 10:

2(x-10)(x 4)

Finally we look back at the sign of 40. It is -,
so the sign of the 4 must be OPPOSITE to the sign
in the first parentheses, so it must be +: 

2(x-10)(x+4)


[If the sign of 40 had been + we would have put 
the SAME sign]

Edwin